OpenStudy (anonymous):

Which conjunction or disjunction equivalent to the open sentence 4 –3|n + 2| ≥ 1? A. n + 2 ≥ 1 and n + 2 ≤ –1 B. n + 2 ≤ 1 and n + 2 ≥ –1 C. n + 2 ≤ 1 and n − 2 ≥ −1 D. n + 2 ≥ 1 or n – 2 ≤ 1

OpenStudy (zzr0ck3r):

\(4-3|n+2|\ge 1\) first subtract \(4\) from both sides \(-3|n+2|\ge -3\) then divide by \(-3\) remembering to flip the sign \(|n+2|\le 1\)

OpenStudy (zzr0ck3r):

with me?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

\(|a|\le b \implies -b\le a \le b \) so we have \(|n+2|\le 1 \implies -1\le n+2\le 1 \implies -3\le n \le -1\) so \(n\le -1\) AND \(-3\le n\)

OpenStudy (zzr0ck3r):

you can actually stop at \(-1\le n+2 \le 1\) this implies \(n+2 \le 1\) and \(-1\le n+2\)